Let ϕ and Ψ be function defined on [-3,5], such that both are continuous on [-3,5], derivable in [-3,5] and ϕ'(x)= Ψ'(x) ∀ x∈]-3,5[. Prove that


ϕ(x)= Ψ(x) +c ∀ x∈ [-3,5] , where c is a real constant

1
Expert's answer
2022-06-06T14:50:42-0400

PROOF

Denote "h(x)=\Phi (x)-\Psi(x)" . The function "\Phi" and "\Psi" are continuous and differentiable on the segment "[-3,5]" , therefore "h" is also differentiable on the "[-3,5]" ."h'(x)=\Phi'(x)-\Psi'(x)=0" for all "x\in""]-3,5[." Let "x_{1}, x_{2}\in ]-3,5[" and "x_{1}<x_{2}" .

By the Mean Value Theorem for the function "h" there exists a point "d\in ]x_{1},x_{2}[" , such that

"h(x_{1})-h(x_{2})=h'(d)\cdot(x_{1}-x_{2})=0" .

Hence, "h(x_{1})-h(x_{2})=0 \Rightarrow h(x_{1})=h(x_{2})" . Since "x_{1}, x_{2}" were any two values of "x\in]-3,5[" this is what it means for a function "h" to be constant on the interval "]-3,5[." So, "h(x)=c" for all "x\in]-3,5[" , where "c\in\R" .

"\lim_{x\rightarrow (-3)^{+}}h(x)= \lim_{x\rightarrow 5^{-}}h(x)=c" . Since "h" is continuous at the points "x=-3" and "x=5" , then "h(-3)=h(5)=c" . Therefore,

"h(x)=c" for all "x\in[-3,5]" "\Rightarrow" "\Phi(x)=\Psi(x)+c" for all "x\in[-3,5]" ,"c\in\R" .